OpenStudy (ayyookyndall):

The x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line? PLEASE HELP!!!!

OpenStudy (imstuck):

If the x intercept is -5, it is -5 when y = 0. If the y intercept is -2, it is -2 when x = 0. So your 2 coordinates to find the equation are (0, -2) and (-5, 0). Can you find the slope of the line if you know those 2 points?

OpenStudy (ayyookyndall):

umm I have options.. y= -5/2x -5 y= 2/5x +2 y= 5/2x -2 y= -2/5x -2

OpenStudy (imstuck):

I am aware you have options. Do you want to know how to do this or are you just looking for answers? I don't need to see the options; I can figure out what they are asking and get the right answer without knowing the options. Do you want to know how to do it or no?

OpenStudy (ayyookyndall):

Yes I am so confused.

OpenStudy (imstuck):

Ok. When you have an x intercept, that means that you replace y with 0 and solve for x. They have already told you that the x intercept is -5. That means that x is -5 and y is 0. The same goes for the y intercept. The y intercept is -2 when x is 0. so you have two points on the line that they are asking you to find the equation for. The 2 points are (-5, 0) and (0, -2).

OpenStudy (imstuck):

When you have two points on a line and you are asked to find the equation, you need to first find the slope of the line. You will eventually need it to get the equation into y=mx+b form, which is the slope-intercept form. Do you know how to find slope from 2 points?

OpenStudy (ayyookyndall):

No. I remember it has something to do with a formula.

OpenStudy (imstuck):

\[m=\frac{ y _{2}-y _{1} }{x _{2} -x _{1}}\]

OpenStudy (imstuck):

The y2 is the y coordinate is in either one of the set of parenthesis. So it could be either set of parenthesis, as long as you start with the same one for both the x and the y.

OpenStudy (imstuck):

Let's start with the point (0, -2) and call that the first set, or x1, y1. The other set will be (-5, 0), or x2 and y2. ok?

OpenStudy (ayyookyndall):

OMG I thought so. so how do I find out what goes for y2 and y1?

OpenStudy (imstuck):

Going from that, we have \[m=\frac{ 0-(-2) }{ -5-0}\]

OpenStudy (imstuck):

Like I said, it doesn't matter which set of coordinates you start with, only that you start with the same one. If you call -5 x1, then 0 will be y1. Get it?

OpenStudy (ayyookyndall):

so it would -2/-5? so y= -2/5x -2

OpenStudy (imstuck):

-2/5 is the slope. Since that is the only choice with a slope of -2/5, then yes, that's the answer. Most times it will not be that easy. But if this is as far as you have gotten in class, and finding the slope is the main point of the lesson and NOT learning how to write the full equation, you're set for now.

OpenStudy (ayyookyndall):

Thank you so much!!!

OpenStudy (imstuck):

Good job. When it gets harder, you'll probably need to come back here for more assistance, but you're welcome for that! Good luck! TY for the medal!

OpenStudy (ayyookyndall):

No problem. One last question though, how did the -2 come in place?

OpenStudy (imstuck):

Ok, this is easy enough. This must be the very beginnings of learning equations, I'm guessing. If you read the problem again, it tells you that you have an x intercept and a y intercept. The x intercept only matters when you are finding slope. The y intercept, however, is the "b" part of the slope-intercept equation: y=mx+b. The m is the slope, which you had to solve for (-2/5) and the y intercept they gave you as -2. So when you solve for the slope, you fill in -2/5 for the m, and the b is substituted with the -2 they gave you in the problem. Do you get that?

OpenStudy (ayyookyndall):

Yes.

OpenStudy (imstuck):

Cool.

OpenStudy (imstuck):

Good job.